OpenStudy (anonymous):

there are 18 gallons of water in a tank. the tank is 3/4 full. how many gallons of water can the tank hold

OpenStudy (anonymous):

18 gallons is 3/4 We can set up a proportion. \(\dfrac{18}{0.75} \rightarrow \dfrac{x}{1}\) Cross multiply and solve: \(18 = 0.75\) Divide 0.7 to both sides, what's 18 / 0.75?

OpenStudy (anonymous):

Sorry that's: \(\dfrac{18}{0.75} = \dfrac{x}{1}\)

OpenStudy (anonymous):

24

OpenStudy (anonymous):

*0.75

OpenStudy (anonymous):

Yes, 24 is correct.

OpenStudy (anonymous):

Welcome to Open Study! You can give medals by clicking 'Best Response'. @Oregon26

OpenStudy (anonymous):

can you help me with another before tou leave please

OpenStudy (anonymous):

i know that ive been a meber

OpenStudy (anonymous):

Sure, just close this one and open a new one.